OpenStudy (anonymous):

find the equation of the line with x-intercept 4 and y-intercept 8 plzz help!!!

OpenStudy (anonymous):

For x-intercept 4 means when y is 0 x is 4 => (4,0) For y-intercept 8 means when x is 0 y is 8 => (0,8) To find the slope, m = (y2 -y1) / (x2 - x1) y2 = 8, y1 = 0 x2 = 0, x1 = 4 m = (8 - 0) / (0 - 4) m = -2 To find the equation of a line, y - y1 = m(x - x1) Taking any one of the coordinates, (0, 8) y - 8 = -2(x - 0) y - 8 = -2x y = -2x + 8 equation of the line is y = -2x + 8

OpenStudy (anonymous):

thank you sooo much !!!

OpenStudy (anonymous):

You welcome :)